Suppose that a one-year forward contract is entered into for a commodity X with a spot price at that time of 1000. Say after 3 months that the spot price for X has increased to 1030. If we assume a force of interest rate of 4%, find the value of the short position in the forward contract after 3 months.
Answer #1

Forward price after 3 months = 1000( 1 + 0.04)3/12

Forward price after 3 months = 1000( 1 + 0.04)0.25

Forward price after 3 months = 1,009.9

Value of short postion = 1009.9 - 1030 = -20.1